OpenStudy (anonymous):

use the quadratic formula: 3x^2-4x+2=0

OpenStudy (anonymous):

\[3x ^{2}-4x+2=0\]

OpenStudy (anonymous):

\[\Large \begin{array}{l} \frac{{ - b \pm \sqrt {{b^2} - 4ac} }}{{2a}}\\ a = 3,b = - 4,c = 2 \end{array}\] Go nuts ;)

OpenStudy (anonymous):

thanks